# Instantaneous Type I blow-up and non-uniqueness of smooth solutions of the Navier–Stokes equations

### v1

@@ 4.1. Error bounds on the velocity
$\begin{split}\|\nabla^{n}\mathcal{I}_{k}^{>t_{k}}\|_{L^{\infty}}&=\left\|\int_{t_{k}}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\nabla^{n}\mathbb{Q}\left(a_{j,k+1}^{2}\theta_{j}\otimes\theta_{j}\right)\,ds\right\|_{L^{\infty}}\\ &\lesssim_{\varepsilon}\int_{t_{k}}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}N_{J_{d},k}^{n+\varepsilon}\,ds\\ &\lesssim\sum_{j}N_{J_{d},k}^{n+\varepsilon}e^{-2N_{j,k+1}^{2}t_{k}},\end{split}$
@@
$\begin{split}\|\nabla^{n}(\bar{R}_{k}(t)-\mathcal{I}_{k}(t))\|_{\infty}&\leq\|\nabla^{n}(\bar{R}_{k}-\tilde{\mathcal{I}}_{k}^{<t_{k}})\|_{\infty}+\|\nabla^{n}(\tilde{\mathcal{I}}_{k}^{<t_{k}}-\mathcal{I}_{k})\|_{\infty}\\ &=\|\nabla^{n}(\tilde{\mathcal{I}}_{k}^{1}+\tilde{\mathcal{I}}_{k}^{2}+\tilde{\mathcal{I}}_{k}^{3}-p\operatorname{Id})\|_{\infty}\\ &\quad+\|\nabla^{n}(\tilde{\mathcal{I}}_{k}^{<t_{k}}-\mathcal{I}_{k}^{<t_{k}})\|_{\infty}+\|\nabla^{n}\mathcal{I}_{k}^{>t_{k}}\|_{\infty}\\ &\lesssim_{\varepsilon,m}\sum_{j}\big(N_{j,k}^{n}e^{-N_{j,k}^{2}t/4}(N_{j,k}^{2}t_{k}+M_{j,k}N_{j,k}^{-1})+N_{j,k}^{n-m}M_{j,k}^{m}\big)\\ &\quad+\sum_{j}\big(M_{j,k}N_{j,k}^{n-1}e^{-N_{j,k}^{2}(t-t_{k})/4}+N_{j,k}^{n-m}M_{j,k}^{m}\big)\\ &\qquad+\sum_{j}N_{J_{d},k}^{n+\varepsilon}e^{-2N_{j,k+1}^{2}t_{k}}+t_{k}N_{J_{d},k}^{n+2+\varepsilon}.\end{split}$
@@
with similar bounds for the remaining terms. Also recall that, for $t_{k}\in[0,t_{k}]$,
@@
Then for any $\varepsilon_{0}>0$ and $0<\alpha<\frac{1}{10}$, we can choose $b$ and $m$ large enough and $\varepsilon$ small enough so that
@@
For $n=0$ the above bound brakes down on the time interval $0,t_{k}]$, and we apply estimates (4.2) and (4.3) instead. ∎
@@
For all $\alpha>0$ sufficiently small, $\epsilon_{0}>0$, and $\bar{n}\in\mathbb{N}$, $A$ can be taken sufficiently large such that the principal part $v$ satisfies
$\|\nabla^{n}(R_{k}(t)-\bar{R}_{k}(t))\|_{L^{\infty}}\leq_{n}\epsilon_{0}N_{1,k+1}^{-\alpha}(t^{-\frac{n}{2}+\alpha}+1),\qquad n=1,2,\dots,\bar{n},$
$\|R_{k}(t)-\bar{R}_{k}(t)\|_{L^{\infty}}\lesssim_{\varepsilon}\varepsilon_{0}N_{1,k+1}^{-\alpha}+\mathbbm{1}_{t\leq t_{k}}N_{J_{d},k}^{\varepsilon},\qquad\text{if}\qquad d=2,$
$\|R_{k}(t)-\bar{R}_{k}(t)\|_{L^{\infty}}\lesssim\varepsilon_{0}N_{1,k+1}^{-\alpha}+\mathbbm{1}_{t\leq t_{k}},\qquad\text{if}\qquad d\geq 3.$
@@
with the first term the main contribution and the other two lower order error. We split
@@
$\begin{split}\mathfrak{E}_{k}^{1}+\mathfrak{E}_{k}^{2}&=\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}t}E_{j,k}\\ &+\sum_{j\neq j^{\prime}}N_{j,k+1}N_{j^{\prime},k+1}e^{-N_{j,k+1}^{2}t-N_{j^{\prime},k+1}^{2}t}E_{j,j^{\prime},k}\\ &+\sum_{j,j^{\prime}}N_{j,k+1}N_{j^{\prime},k+1}e^{-N_{j,k+1}^{2}t-N_{j^{\prime},k+1}t}F_{j,j^{\prime},k},\end{split}$ (4.9)
@@
$F_{j,j^{\prime},k}$ $=\Delta(\psi_{j,k}-N_{j,k}^{-2}a_{j,k}\varphi_{j,k}\theta_{j}\sin(N_{j,k}\eta_{j}\cdot x))\otimes\Delta\psi_{j^{\prime},k}$
$\quad+\Delta(N_{j,k}^{-2}a_{j,k}\varphi_{j,k}\theta_{j}\sin(N_{j,k}\eta_{j}\cdot x))\otimes\Delta(\psi_{j^{\prime},k}-N_{j^{\prime},k}^{-2}a_{j^{\prime},k}\varphi_{j^{\prime},k}\theta_{j^{\prime}}\sin(N_{j^{\prime},k}\eta_{j^{\prime}}\cdot x)).$
@@
$\begin{split}R_{k}=&-\int_{0}^{t}e^{(t-s)\Delta}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\mathbb{Q}(a_{j,k+1}^{2}\theta_{j}\otimes\theta_{j})\,ds\\ &-\int_{0}^{t}e^{(t-s)\Delta}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\mathbb{Q}(a_{j,k+1}^{2}(\varphi^{2}_{j,k+1}\sin^{2}(N_{j,k+1}\eta_{j}\cdot x)-1)\theta_{j}\otimes\theta_{j})\,ds\\ &-\int_{0}^{t}e^{(t-s)\Delta}\sum_{j\neq j^{\prime}}N_{j,k+1}N_{j^{\prime},k+1}e^{-N_{j,k+1}^{2}s-N_{j^{\prime},k+1}^{2}s}\\ &\qquad\quad\cdot\mathbb{Q}(a_{j,k+1}a_{j^{\prime},k+1}\varphi_{j,k+1}\varphi_{j^{\prime},k+1}\sin(N_{j,k+1}\eta_{j}\cdot x)\sin(N_{j^{\prime},k+1}\eta_{j^{\prime}}\cdot x)\theta_{j}\otimes\theta_{j^{\prime}})\,ds\\ &-\int_{0}^{t}e^{(t-s)\Delta}\mathbb{Q}(\mathfrak{E}_{k}^{1}+\mathfrak{E}_{k}^{2})\,ds.\\ =&\mathcal{I}_{k}+\mathcal{J}_{k}^{1}+\mathcal{J}_{k}^{2}+\mathcal{E}_{k}.\end{split}$
@@
$\begin{split}&\quad\operatorname{div}(a_{j,k+1}^{2}(x)(\varphi^{2}_{j,k+1}\sin^{2}(N_{j,k+1}\eta_{j}\cdot x)-1)\theta_{j}\otimes\theta_{j}))\\ &=\nabla a_{j,k+1}^{2}\cdot\theta_{j}(\varphi^{2}_{j,k+1}\sin^{2}(N_{j,k+1}\eta_{j}\cdot x)-1)\theta_{j}.\end{split}$ (4.10)
@@
$\begin{split}&\quad\big\|e^{(t-s)\Delta}\mathcal{R}\mathbb{P}\big(\nabla a_{j,k+1}^{2}\cdot\theta_{j}(\varphi^{2}_{j,k+1}\sin^{2}(N_{j,k+1}\eta_{j}\cdot x)-1)\big)\big\|_{C^{\varepsilon}}\\ &\lesssim_{\varepsilon}M_{j,k+1}^{-1}\|\nabla(a_{j,k+1}^{2})\|_{C^{\varepsilon}}e^{-M_{j,k+1}^{2}(t-s)/4}+M_{j,k+1}^{-m}\|\nabla(a_{j,k+1}^{2})\|_{C^{m,\varepsilon}},\end{split}$ (4.11)
@@
$\begin{split}&\quad\big\|\nabla^{n}e^{(t-s)\Delta}\mathcal{R}\mathbb{P}\big(\nabla a_{j,k+1}^{2}\cdot\theta_{j}(\varphi^{2}_{j,k+1}\sin^{2}(N_{j,k+1}\eta_{j}\cdot x)-1)\big)\big\|_{C^{\varepsilon}}\\ &\lesssim_{\varepsilon}N_{j,k+1}^{n-1}\|\nabla(a_{j,k+1}^{2})\|_{C^{\varepsilon}}e^{-M_{j,k+1}^{2}(t-s)/4}+M_{j,k+1}^{-m+n}\|\nabla(a_{j,k+1}^{2})\|_{C^{m,\varepsilon}}\\ &\lesssim N_{j,k+1}^{n-1}N_{J_{d},k}^{1+\varepsilon}e^{-M_{j,k+1}^{2}(t-s)/4}+M_{j,k+1}^{-m+n}N_{J_{d},k}^{m+1+\varepsilon}\end{split}$ (4.12)
@@
Now we focus on $\mathcal{J}_{k}^{2}$. By Lemma A.1,
$\begin{split}\big\|\nabla^{n}e^{(t-s)\Delta}\mathbb{P}&\nabla\big(a_{j,k+1}a_{j^{\prime},k+1}\varphi_{j,k+1}\varphi_{j^{\prime},k+1}\sin(N_{j,k+1}\eta_{j}\cdot x)\sin(N_{j^{\prime},k+1}\eta_{j^{\prime}}\cdot x)\big)\big\|_{\infty}\\ &\lesssim_{\varepsilon}N_{j\vee j^{\prime},k+1}^{n+1+\varepsilon}e^{-N_{j\vee j^{\prime},k+1}^{2}(t-s)/4}+N_{j\vee j^{\prime},k+1}^{n-m}M_{j\vee j^{\prime},k+1}^{m+1+\varepsilon}.\end{split}$
Then, bounding $L^{\infty}$ by $C^{\varepsilon}$, using that $\mathbb{Q}$ is a Calderón–Zygmund operator, and estimating the heat propagator with (A.2), we have
$\begin{split}&\|\nabla^{n}\mathcal{J}_{k}^{2}(t)\|_{L^{\infty}}\\ &\quad=\Big\|\int_{0}^{t}\nabla^{n}e^{(t-s)\Delta}\sum_{j\neq j^{\prime}}N_{j,k+1}N_{j^{\prime},k+1}e^{-N_{j,k+1}^{2}s-N_{j^{\prime},k+1}^{2}s}\\ &\quad\times\mathbb{Q}(a_{j,k+1}a_{j^{\prime},k+1}\varphi_{j,k+1}\varphi_{j^{\prime},k+1}\sin(N_{j,k+1}\eta_{j}\cdot x)\sin(N_{j^{\prime},k+1}\eta_{j^{\prime}}\cdot x)\theta_{j}\otimes\theta_{j^{\prime}})\,ds\Big\|_{\infty}\\ &\quad\lesssim\int_{0}^{t}\sum_{j^{\prime}<j}N_{j,k+1}N_{j^{\prime},k+1}e^{-N_{j,k+1}^{2}s-N_{j^{\prime},k+1}^{2}s}N_{j,k+1}^{n+\varepsilon}e^{-\frac{1}{4}N_{j,k+1}^{2}(t-s)}\,ds\\ &\quad\quad+\int_{0}^{t}\sum_{j^{\prime}<j}N_{j,k+1}N_{j^{\prime},k+1}e^{-N_{j,k+1}^{2}s-N_{j^{\prime},k+1}^{2}s}N_{j,k+1}^{n-m}M_{j,k+1}^{m+\varepsilon}\,ds\\ &\quad\eqcolon\mathcal{J}_{k}^{2,1}+\mathcal{J}_{k}^{2,2}.\end{split}$ (4.13)
@@
Since $\mathcal{E}^{2}_{k}$ collects the lower order terms from the interactions of different directions $j$ and $j^{\prime}$, $\mathcal{E}^{2}_{k}$ obeys the same estimate of $\mathcal{J}^{2}_{k}$.
@@
$\begin{split}&\quad\Big\|\nabla^{n}\int_{0}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}e^{(t-s)\Delta}\mathbb{Q}E^{1}_{j,k}\,ds\Big\|_{\infty}\\ &\lesssim\int_{0}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}N_{j,k+1}^{-1}\\ &\quad\times\big\|\nabla^{n}e^{(t-s)\Delta}\mathbb{Q}(a_{j,k+1}\varphi_{j,k+1}\nabla(a_{j,k+1}\varphi_{j,k+1})\cdot\eta_{j}\sin(2N_{j,k+1}\eta_{j}\cdot x)\theta_{j}\otimes\theta_{j})\big\|_{\infty}\,ds\\ &\lesssim_{\varepsilon}\int_{0}^{t}\sum_{j}N_{j,k+1}e^{-2N_{j,k+1}^{2}s}N_{j,k+1}^{n+\varepsilon}e^{-2N_{j,k+1}^{2}(t-s)/4}\,ds\\ &\quad+\int_{0}^{t}\sum_{j}N_{j,k+1}e^{-2N_{j,k+1}^{2}s}N_{j,k+1}^{n-m}M_{j,k+1}^{m+1+\epsilon}\,ds\\ &\lesssim\sum_{j}N_{j,k+1}^{n-1+\varepsilon}e^{-\frac{1}{4}N_{j,k+1}^{2}t}(1-e^{-\frac{7}{4}N_{j,k+1}^{2}t}).\end{split}$
Note that $E^{2}_{j,k}$ and $E^{3}_{j,k}$ are of the same order and both contain 0-frequency part, it is thus not necessary to use the heat kernel decay. Instead we estimate them by using $\|e^{t\Delta}f\|_{\infty}\leq\|f\|_{\infty}$, obtaining
$\begin{split}&\quad\Big\|\nabla^{n}\int_{0}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}e^{(t-s)\Delta}\mathbb{Q}(E^{2}_{j,k}+E^{3}_{j,k})\,ds\Big\|_{\infty}\\ &\lesssim_{\varepsilon}\int_{0}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\big\|\nabla^{n}\big(N_{j,k+1}^{-2}(\nabla(a_{j,k+1}\varphi_{j,k+1})\cdot\eta_{j})^{2}\big)\big\|_{C^{\varepsilon}}\,ds\\ &\quad+\int_{0}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\big\|\nabla^{n}\big(N_{j,k+1}^{-2}a_{j,k+1}\varphi_{j,k+1}\Delta(a_{j,k+1}\varphi_{j,k+1})\big)\big\|_{C^{\varepsilon}}\,ds\\ &\lesssim\int_{0}^{t}\sum_{j}e^{-2N_{j,k+1}^{2}s}M_{j,k+1}^{n+2+\varepsilon}\,ds\\ &\lesssim N_{j,k+1}^{-2}M_{j,k+1}^{n+2+\varepsilon}.\end{split}$
Noting that $E^{4}_{j,k}$ obeys the same estimate of $E^{1}_{j,k}$ and $E^{5}_{j,k}$ obeys the same estimate of $E^{2}_{j,k}$, we conclude that
$\|\nabla^{n}\mathcal{E}^{1}_{j,k}\|_{\infty}\lesssim\sum_{j}\big(N_{j,k+1}^{n-1+\varepsilon}e^{-\frac{1}{4}N_{j,k+1}^{2}t}+N_{j,k+1}^{-2}M_{j,k+1}^{n+2+\varepsilon}\big).$

### v2

@@ 5.1. Error bounds on the velocity
$\begin{split}\|\nabla^{n}\mathcal{I}_{k}^{>t_{k}}\|_{L^{\infty}}&\leq\left\|\int_{t_{k}}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\nabla^{n}\mathbb{Q}\left(a_{j,k+1}^{2}\theta_{j}\otimes\theta_{j}\right)\,ds\right\|_{L^{\infty}}\\ &\lesssim_{\varepsilon}\int_{t_{k}}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}N_{J_{d},k}^{n+\varepsilon}\,ds\\ &\lesssim\sum_{j}N_{J_{d},k}^{n+\varepsilon}e^{-2N_{j,k+1}^{2}t_{k}},\end{split}$
@@
$\begin{split}\|\nabla^{n}(\bar{R}_{k}(t)-\mathcal{I}_{k}(t))\|_{\infty}&\leq\|\nabla^{n}(\bar{R}_{k}-\tilde{\mathcal{I}}_{k}^{<t_{k}})\|_{\infty}+\|\nabla^{n}(\tilde{\mathcal{I}}_{k}^{<t_{k}}-\mathcal{I}_{k})\|_{\infty}\\ &=\|\nabla^{n}(\tilde{\mathcal{I}}_{k}^{1}+\tilde{\mathcal{I}}_{k}^{2}+\tilde{\mathcal{I}}_{k}^{3}-(\operatorname{Id}-2\frac{\nabla\otimes\nabla}{\Delta})p)\|_{\infty}\\ &\quad+\|\nabla^{n}(\tilde{\mathcal{I}}_{k}^{<t_{k}}-\mathcal{I}_{k}^{<t_{k}})\|_{\infty}+\|\nabla^{n}\mathcal{I}_{k}^{>t_{k}}\|_{\infty}\\ &\lesssim_{\varepsilon,m}\sum_{j}\big(N_{j,k}^{n}e^{-N_{j,k}^{2}t/4}(N_{j,k}^{2}t_{k}+M_{j,k}N_{j,k}^{-1})+N_{j,k}^{n-m}M_{j,k}^{m}\big)\\ &\qquad+\sum_{j}\big(M_{j,k}N_{j,k}^{n-1}e^{-N_{j,k}^{2}(t-t_{k})/4}+N_{j,k}^{n-m}M_{j,k}^{m}\big)\\ &\qquad+\sum_{j}N_{J_{d},k}^{n+\varepsilon}e^{-2N_{j,k+1}^{2}t_{k}}+t_{k}N_{J_{d},k}^{n+2+\varepsilon}.\end{split}$
@@
with similar bounds for the remaining terms. Also recall that, for $t\in[0,t_{k}]$,
@@
Then for any $\varepsilon_{0}>0$ and $0<\alpha<\frac{1}{10}$, we can choose $A$, $b$, and $m$ large enough and $\varepsilon$ small enough so that
@@
For $n=0$ the above bound breaks down on the time interval $0,t_{k}]$, and we apply estimates (5.2) and (5.3) instead. ∎
@@
For all $\alpha>0$ sufficiently small, $\epsilon_{0}>0$, and $\bar{n}\in\mathbb{N}$, parameters $A$ and $b$ can be taken sufficiently large such that
$\|\nabla^{n}(R_{k}(t)-\bar{R}_{k}(t))\|_{L^{\infty}}\leq_{n}\epsilon_{0}N_{1,k}^{-\alpha}(t^{-\frac{n}{2}+\alpha}+1),\qquad n=1,2,\dots,\bar{n},$
$\|R_{k}(t)-\bar{R}_{k}(t)\|_{L^{\infty}}\lesssim_{\varepsilon}\varepsilon_{0}N_{1,k}^{-\alpha}+\mathbbm{1}_{t\leq t_{k}}N_{J_{d},k}^{\varepsilon},\qquad\text{if}\qquad d=2,$
$\|R_{k}(t)-\bar{R}_{k}(t)\|_{L^{\infty}}\lesssim\varepsilon_{0}N_{1,k}^{-\alpha}+\mathbbm{1}_{t\leq t_{k}},\qquad\text{if}\qquad d\geq 3.$
@@
with the first term the main contribution and the other three lower order errors. We split
@@
$\begin{split}\mathfrak{E}_{k}^{1}+\mathfrak{E}_{k}^{2}&=\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}t}E_{j,k}\\ &+\sum_{j\neq j^{\prime}}N_{j,k+1}N_{j^{\prime},k+1}e^{-N_{j,k+1}^{2}t-N_{j^{\prime},k+1}^{2}t}E_{j,j^{\prime},k}\\ &+\sum_{j,j^{\prime}}N_{j,k+1}N_{j^{\prime},k+1}e^{-N_{j,k+1}^{2}t-N_{j^{\prime},k+1}^{2}t}F_{j,j^{\prime},k},\end{split}$ (5.9)
@@
$F_{j,j^{\prime},k}$ $=\Delta(\psi_{j,k+1}-N_{j,k+1}^{-2}a_{j,k+!}\varphi_{j,k+1}\theta_{j}\sin(N_{j,k+1}\eta_{j}\cdot x))\otimes\Delta\psi_{j^{\prime},k+1}$
$\quad+\Delta(N_{j,k+1}^{-2}a_{j,k+1}\varphi_{j,k+1}\theta_{j}\sin(N_{j,k+1}\eta_{j}\cdot x))$
$\qquad\otimes\Delta(\psi_{j^{\prime},k+1}-N_{j^{\prime},k+1}^{-2}a_{j^{\prime},k+1}\varphi_{j^{\prime},k+1}\theta_{j^{\prime}}\sin(N_{j^{\prime},k+1}\eta_{j^{\prime}}\cdot x)).$
@@
$\begin{split}R_{k}=&-\int_{0}^{t}e^{(t-s)\Delta}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\mathbb{Q}(a_{j,k+1}^{2}\theta_{j}\otimes\theta_{j})\,ds\\ &-\int_{0}^{t}e^{(t-s)\Delta}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\mathbb{Q}(a_{j,k+1}^{2}(\varphi^{2}_{j,k+1}\sin^{2}(N_{j,k+1}\eta_{j}\cdot x)-1)\theta_{j}\otimes\theta_{j})\,ds\\ &-\int_{0}^{t}e^{(t-s)\Delta}\sum_{j\neq j^{\prime}}N_{j,k+1}N_{j^{\prime},k+1}e^{-N_{j,k+1}^{2}s-N_{j^{\prime},k+1}^{2}s}\\ &\qquad\quad\cdot\mathbb{Q}(a_{j,k+1}a_{j^{\prime},k+1}\varphi_{j,k+1}\varphi_{j^{\prime},k+1}\sin(N_{j,k+1}\eta_{j}\cdot x)\sin(N_{j^{\prime},k+1}\eta_{j^{\prime}}\cdot x)\theta_{j}\otimes\theta_{j^{\prime}})\,ds\\ &-\int_{0}^{t}e^{(t-s)\Delta}\mathbb{Q}(\mathfrak{E}_{k}^{1}+\mathfrak{E}_{k}^{2})\,ds\\ =&\mathcal{I}_{k}+\mathcal{J}_{k}^{1}+\mathcal{J}_{k}^{2}+\mathcal{E}_{k}.\end{split}$
@@
$\operatorname{div}(a_{j,k+1}^{2}(x)(\varphi^{2}_{j,k+1}\sin^{2}(N_{j,k+1}\eta_{j}\cdot x)-1)\theta_{j}\otimes\theta_{j})\\ \quad=\nabla a_{j,k+1}^{2}\cdot\theta_{j}(\varphi^{2}_{j,k+1}\sin^{2}(N_{j,k+1}\eta_{j}\cdot x)-1)\theta_{j}.$ (5.10)
@@
$\big\|e^{(t-s)\Delta}\mathcal{R}\mathbb{P}\big(\nabla a_{j,k+1}^{2}\cdot\theta_{j}(\varphi^{2}_{j,k+1}\sin^{2}(N_{j,k+1}\eta_{j}\cdot x)-1)\big)\big\|_{C^{\varepsilon}}\\ \lesssim_{\varepsilon}M_{j,k+1}^{-1}\|\nabla(a_{j,k+1}^{2})\|_{C^{\varepsilon}}e^{-M_{j,k+1}^{2}(t-s)/4}+M_{j,k+1}^{-m}\|\nabla(a_{j,k+1}^{2})\|_{C^{m,\varepsilon}},$ (5.11)
@@
$\begin{split}&\big\|\nabla^{n}e^{(t-s)\Delta}\mathcal{R}\mathbb{P}\big(\nabla a_{j,k+1}^{2}\cdot\theta_{j}(\varphi^{2}_{j,k+1}\sin^{2}(N_{j,k+1}\eta_{j}\cdot x)-1)\big)\big\|_{C^{\varepsilon}}\\ &\quad\lesssim_{\varepsilon}N_{j,k+1}^{n-1}\|\nabla(a_{j,k+1}^{2})\|_{C^{\varepsilon}}e^{-M_{j,k+1}^{2}(t-s)/4}+M_{j,k+1}^{-m+n}\|\nabla(a_{j,k+1}^{2})\|_{C^{m,\varepsilon}}\\ &\quad\lesssim N_{j,k+1}^{n-1}N_{J_{d},k}^{1+\varepsilon}e^{-M_{j,k+1}^{2}(t-s)/4}+M_{j,k+1}^{-m+n}N_{J_{d},k}^{m+1+\varepsilon}\end{split}$ (5.12)
@@
Now we focus on $\mathcal{J}_{k}^{2}$. Bounding $L^{\infty}$ by $C^{\varepsilon}$, using that $\mathbb{Q}$ is a Calderón–Zygmund operator, and estimating the heat propagator with (A.2), we have
$\begin{split}&\|\nabla^{n}\mathcal{J}_{k}^{2}(t)\|_{L^{\infty}}\\ &\quad=\Big\|\int_{0}^{t}\nabla^{n}e^{(t-s)\Delta}\sum_{j\neq j^{\prime}}N_{j,k+1}N_{j^{\prime},k+1}e^{-N_{j,k+1}^{2}s-N_{j^{\prime},k+1}^{2}s}\\ &\qquad\times\mathbb{Q}(a_{j,k+1}a_{j^{\prime},k+1}\varphi_{j,k+1}\varphi_{j^{\prime},k+1}\sin(N_{j,k+1}\eta_{j}\cdot x)\sin(N_{j^{\prime},k+1}\eta_{j^{\prime}}\cdot x)\theta_{j}\otimes\theta_{j^{\prime}})\,ds\Big\|_{\infty}\\ &\quad\lesssim\int_{0}^{t}\sum_{j^{\prime}<j}N_{j,k+1}N_{j^{\prime},k+1}e^{-N_{j,k+1}^{2}s-N_{j^{\prime},k+1}^{2}s}N_{j,k+1}^{n+\varepsilon}e^{-\frac{1}{4}N_{j,k+1}^{2}(t-s)}\,ds\\ &\qquad+\int_{0}^{t}\sum_{j^{\prime}<j}N_{j,k+1}N_{j^{\prime},k+1}e^{-N_{j,k+1}^{2}s-N_{j^{\prime},k+1}^{2}s}N_{j,k+1}^{n-m}M_{j,k+1}^{m+\varepsilon}\,ds\\ &\quad\eqcolon\mathcal{J}_{k}^{2,1}+\mathcal{J}_{k}^{2,2}.\end{split}$ (5.13)
@@
Since $\mathcal{E}^{2}_{k}$ collects the lower order terms from the interactions of different directions $j$ and $j^{\prime}$, $\mathcal{E}^{2}_{k}$ obeys the same estimate as $\mathcal{J}^{2}_{k}$.
@@
$\begin{split}\Big\|&\nabla^{n}\int_{0}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}e^{(t-s)\Delta}\mathbb{Q}E^{1}_{j,k}\,ds\Big\|_{\infty}\\ &\lesssim\int_{0}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}N_{j,k+1}^{-1}\\ &\quad\times\big\|\nabla^{n}e^{(t-s)\Delta}\mathbb{Q}(a_{j,k+1}\varphi_{j,k+1}\nabla(a_{j,k+1}\varphi_{j,k+1})\cdot\eta_{j}\sin(2N_{j,k+1}\eta_{j}\cdot x)\theta_{j}\otimes\theta_{j})\big\|_{\infty}\,ds\\ &\lesssim_{\varepsilon}\int_{0}^{t}\sum_{j}N_{j,k+1}e^{-2N_{j,k+1}^{2}s}N_{j,k+1}^{n+\varepsilon}e^{-2N_{j,k+1}^{2}(t-s)/4}\,ds\\ &\quad+\int_{0}^{t}\sum_{j}N_{j,k+1}e^{-2N_{j,k+1}^{2}s}N_{j,k+1}^{n-m}M_{j,k+1}^{m+1+\epsilon}\,ds\\ &\lesssim\sum_{j}N_{j,k+1}^{n-1+\varepsilon}e^{-\frac{1}{4}N_{j,k+1}^{2}t}(1-e^{-\frac{7}{4}N_{j,k+1}^{2}t}).\end{split}$
Note that $E^{2}_{j,k}$ and $E^{3}_{j,k}$ are of the same order and both contain a zero-frequency part, it is thus not necessary to use the heat kernel decay. Instead we estimate them by using $\|e^{t\Delta}f\|_{\infty}\leq\|f\|_{\infty}$, obtaining
$\begin{split}\Big\|\nabla^{n}&\int_{0}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}e^{(t-s)\Delta}\mathbb{Q}(E^{2}_{j,k}+E^{3}_{j,k})\,ds\Big\|_{\infty}\\ &\lesssim_{\varepsilon}\int_{0}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\big\|\nabla^{n}\big(N_{j,k+1}^{-2}(\nabla(a_{j,k+1}\varphi_{j,k+1})\cdot\eta_{j})^{2}\big)\big\|_{C^{\varepsilon}}\,ds\\ &\quad+\int_{0}^{t}\sum_{j}N_{j,k+1}^{2}e^{-2N_{j,k+1}^{2}s}\big\|\nabla^{n}\big(N_{j,k+1}^{-2}a_{j,k+1}\varphi_{j,k+1}\Delta(a_{j,k+1}\varphi_{j,k+1})\big)\big\|_{C^{\varepsilon}}\,ds\\ &\lesssim\int_{0}^{t}\sum_{j}e^{-2N_{j,k+1}^{2}s}M_{j,k+1}^{n+2+\varepsilon}\,ds\\ &\lesssim\sum_{j}N_{j,k+1}^{-2}M_{j,k+1}^{n+2+\varepsilon}.\end{split}$
Noting that $E^{4}_{j,k}$ obeys the same estimate as $E^{1}_{j,k}$ and $E^{5}_{j,k}$ obeys the same estimate as $E^{2}_{j,k}$, we conclude that
$\|\nabla^{n}\mathcal{E}^{1}_{k}\|_{\infty}\lesssim\sum_{j}\big(N_{j,k+1}^{n-1+\varepsilon}e^{-\frac{1}{4}N_{j,k+1}^{2}t}+N_{j,k+1}^{-2}M_{j,k+1}^{n+2+\varepsilon}\big).$